Calculate the area under the standard normal curve between these values. (Round your answers to four decimal places.)

(a)    

z = −1.2 and z = 1.2




(b)    

z = −2.4 and z = −1.4
Answer #1

Solution :

Given that,  

Using standard normal table ,

(a)

P( -1.2< z < 1.2)

= P(z < 1.2) - P(z < -1.2)

= 0.8849 - 0.1151

= 0.7698

P( -1.2< z < 1.2) = 0.7698

(b)

P(-2.4 < z < -1.4)

= P(z < -1.4) - P(z <-2.4 )

= 0.0808 - 0.0082

= 0.0726

P(-2.4 < z < -1.4) = 0.0726